OpenStudy (tw101):

Geometry problem that im a bit confused on. Using the law of sines, in triangle ABC, if m

OpenStudy (tw101):

@ganeshie8

OpenStudy (anonymous):

So 

OpenStudy (tw101):

okay, what would I do next? i know the law of sines is Sin A/a=SinB/b=SinC/c, but when i substitute it all together it doesnt make sense to me.

OpenStudy (anonymous):

Just use the Law of sines ratio \[\large \frac{sin(105)}{23}~=~\frac{sin<a}{14}\] Get <a alone

OpenStudy (tw101):

okay so would I do the inverse of sin and multiply both sides by 14?

OpenStudy (anonymous):

Well let's do it one step at a time. First is to yes, multiply both sides by 14. Tell me what you get.

OpenStudy (tw101):

sin (1470)/322

OpenStudy (anonymous):

It should look more like \[\large \frac{14 \times sin(105)}{23}\]

OpenStudy (tw101):

oh okay. and then do I inverse both sides?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

So \[\large <a~=~ asin \large( \normalsize \frac{14 \times sin(105)}{23} \large )\]

OpenStudy (tw101):

okay i got 36.01

OpenStudy (anonymous):

That would be your answer :) good job

OpenStudy (tw101):

thank you so much!

OpenStudy (anonymous):

\[\huge \color{aqua}N\color{fuchsia}o \space \color{lime}P \color{orange}r \color{blue}o \color{maroon}b \color{red}l \color{olive}e \color{purple}m \ddot\smile \]

OpenStudy (tw101):

can i ask one more question?